Find the p -value for the hypothesis test. A random sample of size 50 is taken. The sample has a mean of 375 and a standard deviation of 81. H 0 : µ = 400 H a : µ< 400 The p -value for the hypothesis test is . Your answer should be rounded to 4 decimal places.
Answer #1

Solution :

\mu= 400

\bar x =375

\sigma =81

n = 50

This is the left tailed test .

The null and alternative hypothesis is ,

H0 :  \mu = 400

Ha : \mu     <400

Test statistic = z

= (\bar x - \mu ) / \sigma / \sqrt n

= (375 -400) /81 / \sqrt 50

= −2.18

P(z < −2.18 ) = 0.0146

P-value = 0.0146